OpenStudy (anonymous):

5. What is an equation of the line in slope-intercept from? m=3 and y-intercept is (0, 4) A. y = 4x +3 B. y = -3x +4 C. y = 4x -3 D. y = 3x +4

OpenStudy (anonymous):

The equation of a line is given by y = slope*x + y-intercept, or y = m*x + b. Well, m=3, y-intercept = (0,4), so b=4, so the equation is: D) y = 3*x + 4.

OpenStudy (anonymous):

thanks i got that too :)